OpenStudy (minthia):

simplify the expression 5^2 -4 over 5 + 2

rishavraj (rishavraj):

5^2 - 4 = 5^2 - 2^2

OpenStudy (minthia):

that's not one of my answers

rishavraj (rishavraj):

a^2 - b^2 = (a + b)(a - b)

OpenStudy (anonymous):

\[\frac{ 5^2-4 }{ 5+2 }=\frac{ 25-4 }{ 7 }=\frac{ 21 }{ 7 }\]

OpenStudy (anonymous):

\[21 \div 7=?\]

OpenStudy (minthia):

3

OpenStudy (anonymous):

Then 3 is the simplified form of the expression.

OpenStudy (minthia):

thank you